# NeuVAS: Neural Implicit Surfaces for Variational Shape Modeling

### v1

@@ Abstract.
digital geometry processing, point cloud, explicit surface reconstruction, film sticking, restricted power diagram
copyright: noneccs: Computing methodologies Shape modeling
@@ 1. Introduction
To tackle the challenge of preserving sharp feature curves, we assign a location-dependent weight to the smooth energy term. More specifically, the weight at a surface location depends on its distance to sharp feature curves - the weight is larger when its location is away from the sharp feature curves, and the weight vanishes to zero when its location approaches the sharp feature curve. This strategy facilitates the modeling of shape feature curves because it separates the influences of the smoothness terms from the adjacent surfaces sharing a sharp feature curve, as shown in Figure 3(b). A gallery of surfaces constructed using NeuVAS from various sparse inputs is shown in Figure 1.
@@ 2. Related Work
The VIPSS algorithm (Huang et al., 2019) achieves impressive resilience to sparsity and non-uniformity of point clouds, and has even been demonstrated on unstructured point clouds. However, this robustness is achieved through a global smoothness energy that may overlook sharp surface features. Additionally, the complexity of VIPSS is $O(n^{3})$, and it can only construct point clouds up to 5k points.  (Xu et al., 2023) proposed point cloud orientation techniques and applied them to construct unstructured 3D data. They compute the normal vectors based on global consistency and then use Poisson reconstruction (Kazhdan and Hoppe, 2013) to generate the mesh surface. However, their approach is ill-suited for reconstructing curve networks. On one hand,  (Xu et al., 2023) relies on the global consistency of normal vectors to generate smooth surfaces, which makes it incapable of handling curve networks with sharp features. On the other hand, even with accurately specified normal vectors,  (Xu et al., 2023) struggles to generate a reasonable shape due to significant gaps in regions far from the curve network. Furthermore,  (Xu et al., 2023) exhibits high sensitivity to both the number of points and the shape of the curve network. It is limited to point cloud scales below 10K, rendering it inadequate for representing complex geometric shapes.
@@ 3.1. Interpolating Input Curves
Fitting a neural implicit surface $S$ to data points is a well-studied topic  (Gropp et al., 2020; Sitzmann et al., 2020; Atzmon and Lipman, 2020; Ma et al., 2021; Ben-Shabat et al., 2022). In the following, we provide a brief account of the loss terms used in our implementation, and refer to the details in the existing works, e.g. (Sitzmann et al., 2020). For interpolating the curve sketches, our loss terms are based on the Eikonal condition, the Dirichlet condition, and a regularization term to avoid extraneous zero-level set points.
@@
Let $\mathcal{P}$ denote the set of sample points of all the input curve sketches, where the SDF is supposed to be zero. Let $\mathcal{Q}$ be a bounding region of the input data, which is assumed to be normalized to the range $[-0.5,0.5]^{3}$ by default. By the Eikonal condition, an SDF $f(\boldsymbol{x};\Theta)$ possesses a unit gradient, i.e. $\|{\nabla_{\boldsymbol{q}}}f\|=1$, at any point $\boldsymbol{q}\in\mathcal{Q}$. To constrain our neural function $f(\boldsymbol{x};\Theta)$ to represent an SDF, we use the Eikonal condition to define the loss term
(1) $\mathcal{L}_{E}=\frac{1}{|\mathcal{Q}|}\sum_{\boldsymbol{q}\in\mathcal{Q}}\left|1-\|\nabla f(\boldsymbol{q};\Theta)\|\right|.$
@@ 3.3. Creating Sharp Feature Curves
To make the smoothness energy constraints decay around the feature curves, we introduce a weight function $d^{2}(\boldmath{s},\mathcal{P}_{f})$, where $d(\boldmath{s},\mathcal{P}_{f})$ is the Euclidean distance from a surface point $\boldmath{s}\in S$ to $\mathcal{P}_{f}$, the set of points on the input feature curves. See the inset figure. Then, incorporate this weight function to devise the following modified smoothness energy term:
@@ 4.1. Experiment Setup
Similar to various implicit surface reconstruction methods (Wang et al., 2022; Ben-Shabat et al., 2022; Lipman, 2021), our NeuVAS employs the IGR (Gropp et al., 2020) network architecture, featuring 8 hidden layers, each with 256 units, resulting in a model of 1.86M parameters. The activation function used is softplus. Inputs are normalized to the range $[-0.5,0.5]^{3}$ before being fed into the network.
@@
In our experiments, we utilized the thin-plate energy and determined the weights $\lambda_{\text{E}}=0.1$, $\lambda_{\text{DM}}=100$, and $\lambda_{\text{DNM}}=10$. Empirically, we recommend setting $\lambda_{\text{Smooth}}$ to $5\times 10^{-4}$. In each iteration, we sample $\mathcal{Q}$ of 10K points uniformly inside the bounding box (normalized to $[-0.5,0.5]^{3}$); additionally, a $\mathcal{Q}_{zero}$ of 10K points are uniformly sampled on the zero-level set. Throughout the training phase, we applied the Adam optimizer (Kingma and Ba, 2014) with a default learning rate of $5\times 10^{-5}$ and completed training in 10,000 iterations. The experiments were executed on an NVIDIA GeForce RTX 4090 graphics card equipped with 24GB of video memory along with an Intel(R) Core i9-13900k processor.
@@ A.5. Shape Evolution over Training Iterations
Figure 19. Shape evolution over training iterations. $Iter$ represents the number of iterations. The curve types are set to smooth (top), feature (middle), and mixed (bottom), where only the ear is designated as a feature. The results show that the shape stabilizes around 5K iterations.

### v2

@@ Abstract.
submissionid: 1755journal: TOGcopyright: ccjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 6publicationmonth: 12doi: 10.1145/3763331ccs: Computing methodologies Parametric curve and surface models
@@ 1. Introduction
To tackle the challenge of preserving sharp feature curves, we assign a location-dependent weight to the smooth energy term. More specifically, the weight at a surface location depends on its distance to sharp feature curves - the weight is larger when its location is away from the sharp feature curves, and the weight vanishes to zero when its location approaches the sharp feature curve. This strategy facilitates the modeling of shape feature curves because it separates the influences of the smoothness terms from the adjacent surfaces sharing a sharp feature curve, as shown in Figure 3(b). A gallery of surfaces constructed using NeuVAS from various sparse inputs is shown in Figure 1. Figure 1 includes structured curve networks (purple) as well as more challenging inputs: unstructured and imperfect sketches (blue), and sparse points sampled along curves (green). The latter two are commonly recognized as lower-quality input. The high-quality curves are typically used in CAD modeling, and the lower-quality inputs are more common in sketch-based modeling, especially in immersive VR/AR environments. A key contribution of our method is to handle not only clean structured curves but also disconnected and inaccurate curve collections. To further validate the robustness of our method, we add Gaussian noises to the input curves and show the results in Figure 13.
@@ 2. Related Work
The thin-plate RBF (turk2002modelling) and VIPSS (Huang et al., 2019) algorithm achieves impressive resilience to sparsity and non-uniformity of point clouds, and has even been demonstrated on unstructured point clouds. However, this robustness is achieved through a global smoothness energy that may overlook sharp surface features. Moreover, to avoid degenerate solutions such as a constant zero interpolant, thin-plate RBF requires additional spatial constraints with prescribed signed values. Additionally, both thin-plate RBF and VIPSS have a computational complexity of $O(n^{3})$, which limits their practical application to point clouds of up to around 5k points.  (Xu et al., 2023) proposed point cloud orientation techniques and applied them to construct unstructured 3D data. They compute the normal vectors based on global consistency and then use Poisson reconstruction (Kazhdan and Hoppe, 2013) to generate the mesh surface. However, their approach is ill-suited for reconstructing curve networks. On one hand,  (Xu et al., 2023) relies on the global consistency of normal vectors to generate smooth surfaces, which makes it incapable of handling curve networks with sharp features. On the other hand, even with accurately specified normal vectors,  (Xu et al., 2023) struggles to generate a reasonable shape due to significant gaps in regions far from the curve network. Furthermore,  (Xu et al., 2023) exhibits high sensitivity to both the number of points and the shape of the curve network. It is limited to point cloud scales below 10K, rendering it inadequate for representing complex geometric shapes.
@@ 3.1. Interpolating Input Curves
Fitting a neural implicit surface $S$ to data points is a well-studied topic  (Gropp et al., 2020; Sitzmann et al., 2020; Atzmon and Lipman, 2020; Ma et al., 2021; Ben-Shabat et al., 2022; park2019deepsdf). In the following, we provide a brief account of the loss terms used in our implementation, and refer to the details in the existing works, e.g. (Sitzmann et al., 2020). For interpolating the curve sketches, our loss terms are based on the Eikonal condition, the Dirichlet condition, and a regularization term to avoid extraneous zero-level set points.
@@
We use the Eikonal loss to promote an approximate SDF field. First, it helps regularize the implicit field, leading to cleaner zero-level sets, as discussed in (Gropp et al., 2020). Second, our point sampling strategy relies on having an approximate SDF to ensure distances and gradients are reliable near the surface. Let $\mathcal{P}$ denote the set of sample points of all the input curve sketches, where the SDF is supposed to be zero. Let $\mathcal{Q}$ be a bounding region of the input data, which is assumed to be normalized to the range $0.5,0.5]^{3}$ by default. By the Eikonal condition, an SDF $f(\boldsymbol{x};\Theta)$ possesses a unit gradient, i.e. $\|{\nabla_{\boldsymbol{q}}}f\|=1$, at any point $\boldsymbol{q}\in\mathcal{Q}$. To constrain our neural function $f(\boldsymbol{x};\Theta)$ to represent an SDF, we use the Eikonal condition to define the loss term:
[…]
@@ 3.2. Surface Smoothness Energy
Remark.
Note that we use general curvature formulas instead of simpler formulas specific to SDFs. The reason is that, although our network is trained to approximate an SDF via the Eikonal loss, the resulting field is only an approximate SDF and may have considerable deviations from a true distance field, especially during training. As our focus is on accurately estimating curvature near the zero-level set, we chose to use the general formulas for mean and Gaussian curvature that apply to implicit fields, rather than relying on properties specific to true SDFs.
@@ 3.3. Creating Sharp Feature Curves
To make the smoothness energy constraints decay around the feature curves, we introduce a weight function $d^{2}({s},\mathcal{P}_{f})$, where $d({s},\mathcal{P}_{f})$ is the Euclidean distance from a surface point ${s}\in S$ to $\mathcal{P}_{f}$, the set of points on the input feature curves. See the inset figure. Then, incorporate this weight function to devise the following modified smoothness energy term:
@@ 4.1. Experiment Setup
Similar to various implicit surface reconstruction methods (Wang et al., 2022; Ben-Shabat et al., 2022; Lipman, 2021), our NeuVAS employs the IGR (Gropp et al., 2020) network architecture, featuring 8 hidden layers, each with 256 units, resulting in a model of 1.86M parameters. The activation function used is softplus. Inputs are normalized to the range $0.5,0.5]^{3}$ before being fed into the network.
@@
In our experiments, we utilized the thin-plate energy and determined the weights $\lambda_{\text{E}}=0.1$, $\lambda_{\text{DM}}=100$, and $\lambda_{\text{DNM}}=10$. Empirically, we recommend setting $\lambda_{\text{Smooth}}$ to $5\times 10^{-4}$. In each iteration, we sample $\mathcal{Q}$ of 10K points uniformly inside the bounding box (normalized to $0.5,0.5]^{3}$); additionally, a $\mathcal{Q}_{zero}$ of 10K points are uniformly sampled on the zero-level set. Throughout the training phase, we applied the Adam optimizer (Kingma and Ba, 2014) with a default learning rate of $5\times 10^{-5}$ and completed training in 10,000 iterations. For visualization, meshes are extracted from the zero-level set using the Marching Cubes algorithm at a consistent resolution of $512^{3}$ for all methods under comparison. The experiments were executed on an NVIDIA GeForce RTX 4090 graphics card equipped with 24GB of video memory along with an Intel(R) Core i9-13900k processor.
@@ 5. Limitation
Sharpness
Due to the inherent limitations of the SDF representation, our method, despite achieving a high degree of sharpness, is unable to attain perfectly sharp features. This behavior is illustrated in Figure 20.
Euclidean distance weight
In extreme cases, the shape may contain two very close, nearly parallel surfaces. These surfaces are disconnected from each other, one is defined by a smooth curve, while the other is defined by a sharp curve. Under such conditions, interactions between the two surfaces can occur, leading to slight irregularities or a reduction in smoothness.
@@ 6. Conclusion
Acknowledgements.
The authors would like to thank the anonymous reviewers for their valuable comments and suggestions. This work was supported by the National Key Research and Development Program of china (2024YFB3309500), the National Natural Science Foundation of China (62402295, 62172257, U23A20312, 62272277), the Natural Science Foundation of Shandong Province (ZR2024QF087, ZR2024ZD12), the Innovation and Technology Commission of the HKSAR Government under the InnoHK initiative and the Hong Kong RGC (Ref. T45-205/21-N), the Joint Fund of the National Natural Science Foundation of China (U22A2033, U24A20219), the National Key R&D Program of China (2022YFB3303200), and the Natural Sciences and Engineering Research Council of Canada (RGPIN-2024-03981).
@@ A.5. Shape Evolution over Training Iterations
Figure 19. Shape evolution over training iterations. IterIter represents the number of iterations. The curve types are set to smooth (top), feature (middle), and mixed (bottom), where only the ear is designated as a feature. The results show that the shape stabilizes around 5K iterations.
@@
A.6. Sharpness
(a) Input (b) $Res:128^{3}$ (c) $Res:256^{3}$ (d) $Res:512^{3}$ (e) $Res:1024^{3}$
Figure 20. Marching Cubes at multiple resolutions are used to demonstrate the sharpness.
We employ Marching Cubes at multiple resolutions to visualize the sharpness of the constructed surfaces, as shown in Figure 20. The resolution of the Marching Cubes is progressively increased to clearly illustrate the improvement in sharp feature, accompanied by 2D SDF slices for further clarity. At a resolution of $1024^{3}$, the constructed surface closely approximates the true SDF. Although perfect sharpness is not achieved, the results exhibit a high degree of geometric fidelity and are visually near-sharp.
A.7. Additional datasets
Figure 21. A gallery of our results on  (Yu et al., 2022) dataset.
We present results on the  (Yu et al., 2022) dataset in Figure 21 to further demonstrate the generality and robustness of our approach. Since our method has difficulty handling open surfaces, models containing such geometries were excluded from the evaluation.